### v1

# On the spectral radius of the ratio of Girko matrices

Abstract.
Girko matrices have independent and identically distributed entries of mean zero and unit variance. In this note, we consider the random matrix model formed by the ratio of two independent Girko matrices, its entries are dependent and heavy-tailed. Our main message is that divided by the square root of the dimension, the spectral radius of the ratio converges in distribution, when the dimension tends to infinity, to a universal heavy-tailed distribution. We provide a mathematical proof of this high-dimensional phenomenon, under a fourth moment matching with a Gaussian case known as the complex Ginibre ensemble. In this Gaussian case, the model is known as the spherical ensemble, and its spectrum is a determinantal planar Coulomb gas. Its image by the inverse stereographic projection is a rotationally invariant gas on the two-sphere. A crucial observation is the invariance in law of the model under inversion, related to its spherical symmetry, and that makes, in a sense, edge and bulk equivalent. Our approach involves Girko Hermitization, local law estimates for Wigner matrices, lower bound estimates on the smallest singular value, and convergence of kernels of determinantal point processes. The universality of the high-dimensional fluctuation of the spectral radius of the ratio of Girko matrices turns out to be remarkably more accessible mathematically than for a single Girko matrix!
Key words and phrases:
Random matrix; Spherical Model; Spectral Radius; Logarithmic Potential; Fluctuation; Universality
2020 Mathematics Subject Classification:
60B20 - Random matrices (probabilistic aspects); 15A60 - Norms of matrices, numerical range, applications of functional analysis to matrix theory; 60F05 - Central limit and other weak theorems.
1. Introduction
The asymptotic analysis in high dimension of the spectral radius of non-normal random matrices is a delicate subject, which goes back at least to Cohen and Newman (1984); Geman (1986); Bai and Yin (1986); Rider (2003). There has been recent progress in the case of Girko random matrices, see for instance Bordenave et al. (2022); Cipolloni et al. (2024a, b); Byun and Forrester (2025) and references therein. Girko matrices have independent square-integrable entries. This note addresses the case of the ratio of two independent Girko matrices, a model of non-normal random matrices with dependent entries. For this model, we prove the universality of the asymptotic behavior of the spectral radius in high dimension. This model, in the special case of Ginibre matrices, known as the spherical ensemble, becomes integrable and biunitary invariant, and the law of its spectrum is a planar heavy-tailed Coulomb gas, with a determinantal structure. The spherical ensemble is a matrix analogue and extension of the ratio of independent real normal random variables, and can be then seen as a sort of matrix Cauchy law.
The eigenvalues of an $n\times n$ complex matrix $A=(A_{jk}:1\leq j,k\leq n)\in\mathcal{M}_{n}(\mathbb{C})$ are the roots in $\mathbb{C}$ of its characteristic polynomial $\det(A-z\mathrm{Id}_{n})\in\mathbb{C}[z]$. They form a multiset of cardinal $n$ called the spectrum of $A$, denoted $\mathrm{spec}(A)$. The empirical spectral distribution, spectral radius, and reciprocal or inner spectral radius of $A$ are [^1] Obviously when $A$ is invertible, then $\rho_{\min}(A)=\rho_{\max}(A^{-1})^{-1}$, a formula that remains valid in an extended sense when $A$ is singular.
$\mu_{A}=\frac{1}{n}\sum_{\lambda\in\mathrm{spec}(A)}\delta_{\lambda},\qquad\rho_{\max}(A)=\max_{\lambda\in\mathrm{spec}(A)}|\lambda|,\quad\text{and}\quad\rho_{\min}(A)=\min_{\lambda\in\mathrm{spec}(A)}|\lambda|.$ (1.1)
1.1. Main results
We say that an $n\times n$ random matrix $A=(A_{jk}:1\leq j,k\leq n)$ is a Girko matrix when its entries are independent and identically distributed (i.i.d.) in $\mathbb{C}$ with $\mathbb{E}(A_{11})=0$ and $\mathbb{E}(|A_{11}|^{2})=1$. We say that $A$ is a complex Ginibre matrix in the special case where $A_{11}\sim\mathcal{N}_{\mathbb{C}}(0,\frac{1}{2}\mathrm{Id}_{2})$, which is equivalent to state that $A$ has density in $\mathbb{C}^{n\times n}$ proportional to $A\in\mathcal{M}_{n}(\mathbb{C})\mapsto\exp(-\mathrm{Tr}(AA^{*}))$. To distinguish, we denote such a Ginibre matrix by $A^{\mathrm{Gin}}$, and we always write $\mathrm{Gin}$ as a superscript to indicate this special case.
We consider the following assumptions or conditions on an $n\times n$ Girko matrix $A$:
  1. (C1)
    the law of $A_{11}$ has a bounded density function $\varphi$ with respect to the Lebesgue measure on $\mathbb{C}$, and
    $\|\varphi\|_{\infty}\leq D_{0}\quad\text{ for some constant $D_{0}>0$.}$
  2. (C2)
    Fourth moment matching condition. There exists a small constant $c_{0}>0$ such that
    $\mathcal{E}_{1}=\mathcal{E}_{2}=0,\quad\qquad\mathcal{E}_{3}\leq n^{-\frac{1}{2}-c_{0}},\quad\qquad\mathcal{E}_{4}\leq n^{-c_{0}},$
    where ($k$-th moment deviation of $A_{11}$ from standard complex Gaussian)
    $\mathcal{E}_{k}=\max_{\begin{subarray}{c}a,b\in{\mathbb{N}}\\ a+b=k\end{subarray}}\left|\mathbb{E}[(\operatorname{Re}A_{11})^{a}(\operatorname{Im}A_{11})^{b}]-\mathbb{E}[(\operatorname{Re}A^{\mathrm{Gin}}_{11})^{a}(\operatorname{Im}A^{\mathrm{Gin}}_{11})^{b}]\right|.$
  3. (C3)
    Finite moment condition. For each $k\geq 1$, there exists constants $D_{k}>0$ (independent of $n$) such that
    $\mathbb{E}[|A_{11}|^{k}]<D_{k}.$
Condition (C1) implies that the $n\times n$ complex matrix $A$ is a.s. invertible for all $n$ since the algebraic hypersurface $\{\det=0\}$ has zero Lebesgue measure. Moreover, the uniform bound on the density allows a precise lower bound on the smallest singular value, see e.g., Jain et al. (2021). More generally, we refer to Tikhomirov (2023) for the invertibility of random matrices beyond the density case.
Condition (C2) states that the moments of $A_{11}$ almost match those of $A^{\mathrm{Gin}}_{11}$ up to the fourth order.
Condition (C3) is assumed for technical reasons to simplify a proof.
Example 1.1 (Moment matching with the standard complex Gaussian distribution).
An example of a probability measure $\nu$ on $\mathbb{C}\equiv\mathbb{R}^{2}$ with moment matching up to order $k\geq 1$ with $\mu=\mathcal{N}(0,\frac{1}{2}\mathrm{Id}_{2})$ is
$\mathrm{d}\nu=(1+f)\mathrm{d}\mu$ (1.2)
where $f\in L^{2}(\mu)$ is such that $f\neq 0$, $f\geq-1$, and $f\perp\mathbb{R}_{k}[X,Y]$ where $\mathbb{R}_{k}[X,Y]$ is the set of bivariate polynomials with total degree less than or equal to $k$, the span of $X^{a}Y^{b}$ with $a,b\in\mathbb{N}$, $a+b\leq k$. Indeed, for all $P\in\mathbb{R}_{k}[X,Y]$,
$\int P\mathrm{d}\nu-\int P\mathrm{d}\mu=\langle P,f\rangle_{L^{2}(\mu)}=0.$ (1.3)
The special case $P=1$ states that $\nu$ is a probability measure. However, this construction has no reason to produce a product measure. Also, we may prefer to take alternatively
$\mathrm{d}\nu=g\mathrm{d}\mu\quad\text{where}\quad g(z)=g(x+\mathrm{i}y)=(1+f_{1}(x))(1+f_{2}(y))$ (1.4)
where $f_{1},f_{2}:\mathbb{R}\to\mathbb{R}$ are in $L^{2}(\mathcal{N}(0,\frac{1}{2}))$ and satisfy $f_{i}\neq 0$, $f_{i}\geq-1$, and $f_{i}\perp\mathbb{R}_{k}[X]$ where $\mathbb{R}_{k}[X]$ is the span of univariate polynomials with total degree less than or equal to $k$. A basic example is given by $f_{1}=f_{2}=c^{-1}H_{m}$ where $H_{m}$ is the Hermite polynomial, orthogonal with respect to $\mathcal{N}(0,\frac{1}{2})$, of even degree $m>k$, and $c=|\inf_{\mathbb{R}}H_{m}|>0$. In this case $\nu$ has a bounded density with respect to the Lebesgue measure on $\mathbb{C}\equiv\mathbb{R}^{2}$. The moment matching up to order $k$ between $\mu$ and $\nu$ comes from $\mu=\mathcal{N}(0,\frac{1}{2})^{\otimes 2}$. In another direction, by using the Gauss – Hermite quadrature, it is possible to construct a finite discrete distribution that matches the moments of the Gaussian up to an arbitrary fixed order.
Theorem 1.2 (From ratio of Girko matrices to infinite Ginibre).
Let $M=AB^{-1}$ where $A$ and $B$ are independent $n\times n$ Girko matrices, not necessarily of the same law, both satisfying (C1)(C3). Let $\mathrm{Gin}_{\infty}$ be the infinite Ginibre determinantal point process on $\mathbb{C}$, with kernel $K(z,w)=\sqrt{\gamma(z)\gamma(w)}\mathrm{e}^{z\overline{w}}$ where $\gamma(z)=\frac{1}{\pi}\mathrm{e}^{-|z|^{2}}$. Then, for all fixed $\lambda_{0}\in\mathbb{C}$, for the topology of continuous and compactly supported test functions,
$n\mu_{\sqrt{n}(1+|\lambda_{0}|^{2})^{-1}(M-\lambda_{0}\mathrm{Id})}=\sum_{\lambda\in\mathrm{spec}(M)}\delta_{\sqrt{n}(1+|\lambda_{0}|^{2})^{-1}(\lambda-\lambda_{0})}\xrightarrow[n\to\infty]{\mathrm{d}}\mathrm{Gin}_{\infty}.$
Theorem 1.3 (Spectral radius).
Let $M$ be as in Theorem 1.2, and ${(\gamma_{k})}_{k\geq 1}$ independent with $\gamma_{k}\sim\mathrm{Gamma}(k,1)$. Then
$\frac{\rho_{\max}(M)}{\sqrt{n}}\xrightarrow[n\to\infty]{\mathrm{d}}\mathrm{Law}\Bigr(\frac{1}{\sqrt{\min_{k\geq 1}\gamma_{k}}}\Bigr)\quad\text{and}\quad\sqrt{n}\rho_{\min}(M)\xrightarrow[n\to\infty]{\mathrm{d}}\mathrm{Law}\bigr(\sqrt{\min_{k\geq 1}\gamma_{k}}\bigr).$ (1.5)
Note that the laws of $R_{0}=\sqrt{\min_{k\geq 1}\gamma_{k}}$ and $R_{\infty}=\displaystyle\frac{1}{R_{0}}$ have respective cdf given for all $x>0$ by
$\mathbb{P}(R_{0}\leq x)=1-\prod_{k=1}^{\infty}(1-F_{k})(x^{2})\quad\text{and}\quad\mathbb{P}(R_{\infty}\leq x)=\prod_{k=1}^{\infty}(1-F_{k})(x^{-2})$ (1.6)
where $F_{k}(x)=\mathbb{P}(\gamma_{k}<x)=\mathbb{P}(\pi_{x}<k)=\mathrm{e}^{-x}\sum_{j=0}^{k-1}\frac{x^{j}}{j!}$ with $\pi_{x}\sim\mathrm{Poisson}(x)$. In particular $R_{\infty}$ is heavy-tailed as
$\mathbb{P}(R_{\infty}>x)\underset{x\to+\infty}{\sim}\frac{1}{x^{2}}.$ (1.7)
Since ${(\xi_{k})}_{k\geq 1}$ are not i.i.d., we are beyond classical extreme value theory : $R_{0}$ is not Weibull, and $R_{\infty}$ is not Fréchet.
Theorem 1.3 establishes, for the first time and to some extent, the universality of the high-dimensional behavior of the spectral radius of the spherical ensemble, studied by Jiang and Qi (2017) when $A$ and $B$ are complex Ginibre matrices, see Section 2. Theorem 1.3 is an analogue for the spherical ensemble of the universality of the Gumbel fluctuation for the complex Ginibre ensemble obtained by Cipolloni et al. (2024b).
The proofs of Theorem 1.2 and Theorem 1.3 are given in Section 3. The strategy is as follows. By using the invariance of the model by inversion, we transform the behavior of the dilated spectral radius into the local behavior near the origin, more precisely a local gap probability at the origin. We then use the universality of the local behavior near the origin, stated as a replacement principle in Theorem 1.4 below. This reduces the problem to the local gap probability of the spherical ensemble, which in turn reduces to the infinite Ginibre ensemble by using kernel convergence of determinantal point processes. Finally, the local gap probability for the infinite Ginibre ensemble follows from the Kostlan observation of the moduli of the particles of this determinantal point process.
The spherical or Forrester – Krishnapur ensemble or model of dimension $n$ is the random matrix defined by
$M^{\mathrm{Gin}}=A^{\mathrm{Gin}}(B^{\mathrm{Gin}})^{-1}$ (1.8)
where $A^{\mathrm{Gin}}$ and $B^{\mathrm{Gin}}$ are independent $n\times n$ complex Ginibre matrices with i.i.d. $\mathcal{N}_{\mathbb{C}}(0,\frac{1}{2}\mathrm{Id}_{2})$ entries.
Theorem 1.4 (Replacement principle for local eigenvalue statistics near a point).
Let $M=AB^{-1}$ be as in Theorem 1.2 and $M^{\mathrm{Gin}}$ as in (1.8). Let $\mu_{M}$ and $\mu_{M^{\mathrm{Gin}}}$ be the empirical spectral measures as in (1.1). Then the following replacement principle on scale $\sqrt{n}$ occurs: for all $z_{0}\in\mathbb{C}$ and all compactly supported $\mathcal{C}^{\infty}$ functions $f:\mathbb{C}\to\mathbb{R}$ and $F:\mathbb{R}\to\mathbb{R}$,
$\lim_{n\rightarrow\infty}\Bigr(\mathbb{E}\Bigr[F\Bigr(\int f(\sqrt{n}(z-z_{0}))\mathrm{d}\mu_{M}(z)\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f(\sqrt{n}(z-z_{0}))\mathrm{d}\mu_{M^{\mathrm{Gin}}}(z)\Bigr)\Bigr]\Bigr)=0.$
Theorem 1.4 is a simplified version of a more technical result stated in Theorem 4.1 and proved in Section 4. It can be seen as a fourth moment theorem for the spherical ensemble. The proof involves Ornstein – Uhlenbeck interpolation between Gaussian and non-Gaussian, a cumulant expansion exploiting the fourth moment matching (C2), a local law for Wigner matrices and its consequence in terms of rigidity, and a precise lower bound on the smallest singular value. The continuous interpolation that we use, based on the Ornstein – Uhlenbeck diffusion, differs from the original Lindeberg-type entry by entry replacement used by Tao and Vu (2015) in their original fourth moment theorem.
Remark 1.5 (Real case).
The replacement principle stated in Theorem 1.4 also applies to the real case where $M=AB^{-1}$ with $A,B$ being real Girko matrices, and $A^{\mathrm{Gin}}$ and $B^{\mathrm{Gin}}$ being real Ginibre matrices. The proof is essentially the same with slight modifications. For the joint eigenvalue distribution of the real spherical ensemble, we refer to Forrester and Mays (2012).
1.2. Comments and open problems
1.2.1. Model.
We could use the alternative definitions $M=A^{-1}B$ instead of $M=AB^{-1}$. Indeed $XY$ and $YX$ have same spectrum thanks to the Sylvester identity $\det(XY-z\mathrm{Id})=\det(YX-z\mathrm{Id})$.
1.2.2. First order universality.
The universality of the limiting spectral distribution of the model was proved in Bordenave (2011), using logarithmic potential, Girko Hermitization, and Tao and Vu replacement principle. The limit can be interpreted by using the notions of freeness and Brown measure from Free Probability Theory.
1.2.3. Inner radius of Girko matrices.
The analogue of Theorem 1.2 for a single Girko matrix AA follows from (Maltsev and Osman, 2024, Th. 2.1 and 2.2), under a second moment matching assumption. This leads to the universality of the fluctuation of the inner spectral radius, which meets the gap at the origin of $\mathrm{Gin}_{\infty}$, namely
$\sqrt{n}\rho_{\min}(A)\xrightarrow[n\to\infty]{\mathrm{d}}\mathrm{Law}\Bigr(\sqrt{\min_{k\geq 1}\xi_{k}}\Bigr)\quad\text{where}\quad\text{${(\xi_{k})}_{k\geq 1}$ are independent with $\xi_{k}\sim\mathrm{Gamma}(k,1)$.}$ (1.9)
1.2.4. Moments assumptions.
The fourth moment matching in (C2) helps to reduce the complexity of the proof of Theorem 4.1, essentially the bulk universality for Girko matrices, see for instance Tao and Vu (2015). Theorem 1.3 and Theorem 4.1 are expected to remain available with the sole second moment matching.
1.2.5. Characteristic polynomials
It is natural to ask about the asymptotic analysis of the characteristic polynomial of $AB^{-1}$, in the spirit of Bordenave et al. (2022), with a random analytic object as a limit. Taking the modulus and the logarithm would then recover the CLT for the log-potential. This seems to be open even for the spherical model.
1.2.6. Ratio of other models
We could ask about the behavior of the spectral radius of $AB^{-1}$ if $A$ and $B$ are independent real or quaternion Girko matrices, and, in another direction, real/complex/quaternion Wigner matrices.
1.3. Conventions and notation
The convergence in distribution of a sequence of random variables is the weak convergence of their laws, with respect to continuous and bounded test functions, towards a probability distribution. We denote it by a roman letter $\mathrm{d}$ over $\to$”. Similarly, we use $\mathrm{d}$ over “==” for the equality in distribution.
For integers $k,l\in{\mathbb{N}}$, with $k<l$, we write $k,l]=\{k,\dots,l\}$ and $[k]=[1,k]$.
We set $\mathrm{d}^{2}z=\frac{1}{2}\mathrm{i}(\mathrm{d}z\wedge\mathrm{d}\overline{z})$ for the two dimensional area form on $\mathbb{C}$}, matching the Lebesgue measure $\mathrm{d}x\mathrm{d}y$ on $\mathbb{R}^{2}$.
The two-dimensional Laplace operator is $\Delta=\Delta_{z}=4\partial_{z}\partial_{\bar{z}}=\partial^{2}_{\operatorname{Re}z}+\partial^{2}_{\operatorname{Im}z}$ denotes the Laplace operator.
For any vector $\mathbf{x}\in{\mathbb{C}}^{d}$ and matrix $A\in{\mathbb{C}}^{d\times d}$, $d\in{\mathbb{N}}$, we use $\lVert\mathbf{x}\rVert$ and $\lVert A\rVert$ indicate the usual Euclidean norm and the corresponding induced matrix norm (also known as operator norm). Throughout the paper $c,C>0$ denote small and large constants, respectively, which may change from line to line. For positive quantities $f,g$ we write $f\lesssim g$ and $f\sim g$ if $f\leq Cg$ or $cg\leq f\leq Cg$, respectively, for some constants $c,C>0$ independent of $n$. Furthermore, for nn–dependent positive quantities $f_{n},g_{n}$ we use the notation $f_{n}\ll g_{n}$ to denote that $\lim_{n\to\infty}(f_{n}/g_{n})=0$.
We will often use the concept of “with very high probability” for an $n$-dependent event, meaning that for any fixed $D>0$ the probability of the event is bigger than $1-n^{-D}$ when $n\geq n_{0}(D)$. We recall the standard notion of stochastic domination: given two families of non-negative random variables
$X=\left(X^{(n)}(u)\,:\,n\in{\mathbb{N}},u\in U^{(n)}\right)\quad\text{and}\quad Y=\left(Y^{(n)}(u)\,:\,n\in{\mathbb{N}},u\in U^{(n)}\right)$
indexed by $n$ (and possibly some parameter $u$ in some parameter space $U^{(n)}$), we say that $X$ is stochastically dominated by $Y$, if for any small $\xi$ and large $D>0$ we have
$\sup_{u\in U^{(n)}}\mathbb{P}\left[X^{(n)}(u)>n^{\xi}Y^{(n)}(u)\right]\leq n^{-D}$ (1.10)
for large enough $n\geq n_{0}(\xi,D)$. In this case we use the notation $X\prec Y$ or $X=O_{\prec}(Y)$ and we say that $X(u)\prec Y(u)$ holds uniformly in $u\in U^{(n)}$. Properties of stochastic domination can be found in (Erdős and Yau, 2017, Proposition 6.5). We often use the notation $\prec$ also for deterministic quantities, then (1.10) holds with probability one.
Acknowledgments
The authors were supported by the Swedish Research Council under grant no. 2021-06594 while in residence at Institut Mittag-Leffler in Djursholm, Sweden during the fall semester of 2024, Program ”Random Matrices and Scaling Limits” organized by Maurice Duits, Kurt Johansson, Gaultier Lambert, and Kevin Schnelli, KTH Royal Institute of Technology.
Figure 1. Sample of the spherical ensemble and its planar stereographic projection.
[…]
Figure 2. The point $T(P)\in\mathbb{C}$ is the image of $P\in\mathbb{S}^{2}\setminus\{N\}$ by the stereographic projection or transform $T$ with respect of the north pole NN, see (2.7). See also Needham (2023). The image of $N$ is $\infty$. For gases, the gap around NN on $\mathbb{S}^{2}$ is in correspondence with the gap around $\infty$ on $\mathbb{C}\cup\{\infty\}$.
Figure 3. Histograms illustrating the universality of the high dimensional behavior, for three choices of entries law : symmetric complex Bernoulli or Rademacher, isotropic complex Gaussian, and symmetric complex discrete heavy tailed (same tail as Zipf or Pareto, $\alpha=4$). For the Bernoulli and Zipf cases, the matrices are also conditioned to be nonsingular by using rejection sampling.
2. The spherical ensemble
This section gathers useful properties of the spherical ensemble, in potential theory, geometry, and probability.
2.1. Coulomb gas and determinantal structure
The spherical or Forrester – Krishnapur random matrix ensemble (1.8) is considered for instance in Hough et al. (2009); Forrester (2010). The law of $M^{\mathrm{Gin}}$ inherits the biunitary invariance : invariance by multiplication from the left and from the right by deterministic unitary matrices. The spectrum of $M^{\mathrm{Gin}}$ forms a Coulomb gas on $\mathbb{C}$ with density [^2] By conjugacy with an independent Haar unitary matrix, this can also be seen as the spectrum of a random normal matrix model, which should not be confused with the random non-normal matrix model obtained by taking the ratio of two independent complex Ginibre matrices. $\varphi_{n}:\mathbb{C}^{n}\mapsto(0,+\infty)$ given by
$\varphi_{n}(z_{1},\ldots,z_{n})=c_{n}\frac{\prod_{i<j}|z_{i}-z_{j}|^{2}}{\prod_{i=1}^{n}(1+|z_{i}|^{2})^{n+1}}=c_{n}\mathrm{e}^{-(n+1)\sum_{i=1}^{n}Q(|z_{i}|)}\prod_{i<j}|z_{i}-z_{j}|^{2},$ (2.1)
where $c_{n}$ is a normalizing constant. This is a planar Coulomb gas with rotationally invariant potential
$Q(|z|)=\log(1+|z|^{2}).$ (2.2)
This is also a determinantal point process of $n$ particles on $\mathbb{C}$ endowed with the Lebesgue measure with kernel [^3] Contrary to Hough et al. (2009), we use here as a background the uniform distribution on the sphere in stereographic coordinates.
$(z,w)\in\mathbb{C}^{2}\mapsto K_{n}(z,w)=\sqrt{\kappa(z)\kappa(w)}n\Bigr(\frac{1+z\overline{w}}{\sqrt{(1+|z|^{2})(1+|w|^{2})}}\Bigr)^{n-1}\quad\text{where}\quad\kappa(z)=\frac{1}{\pi(1+|z|^{2})^{2}}.$ (2.3)
This means that the joint density in (2.1) is given by
$\varphi_{n}(z_{1},\ldots,z_{n})=\det\bigr[K_{n}(z_{i},z_{j})\bigr]_{1\leq i,j\leq n}.$ (2.4)
The function $\kappa$ in (2.3) is a density on $\mathbb{C}$ with respect to the Lebesgue measure. The probability measure
$\mathrm{d}\nu(z)=\kappa(z)\mathrm{d}^{2}z$ (2.5)
is known as a bivariate Student t distribution in Statistics and as a Barenblatt profile in Analysis of PDE (fast diffusion equation [^4] See for instance Chafaï et al. (2025) and references therein.). It is heavy tailed. It is the image by the (north pole) stereographic projection
$T:\mathbb{S}^{2}\subset\mathbb{R}^{3}\to\mathbb{C}\cup\{\infty\}$ (2.6)
of the uniform probability measure on the sphere $\mathbb{S}^{2}$. More precisely, for all $x=(x_{1},x_{2},x_{3})\in\mathbb{S}^{2}\setminus\{(0,0,1)\}$ and $z\in\mathbb{C}$,
$T(x)=\frac{x_{1}+\mathrm{i}x_{2}}{1-x_{3}}\quad\text{and}\quad T^{-1}(z)=\frac{(2\operatorname{Re}z,2\operatorname{Im}z,|z|^{2}-1)}{|z|^{2}+1}.$ (2.7)
In other words, this measure is the uniform probability distribution on the sphere written in stereographic coordinates. Furthermore, the image of the spherical ensemble by $T^{-1}$ is the gas on $\mathbb{S}^{2}$ with density with respect to the uniform measure given, up to a multiplicative normalizing constant, by
$(x_{1},\ldots,x_{n})\in(\mathbb{S}^{2})^{n}\mapsto\prod_{i<j}\|x_{i}-x_{j}\|^{2}_{\mathbb{R}^{3}},$ (2.8)
hence the name of the ensemble. This can also be seen as a two-dimensional analogue of the circular unitary ensemble (CUE). For all these aspects, we refer to (Hough et al., 2009, Sec. 4.3.8). This point process on the $\mathbb{S}^{2}$ is invariant by isometries, which allows us to study the behavior near $\infty$ by dealing with the behavior near 0. This symmetry is a key ingredient of our proof of Theorem 1.3.
When $R$ runs over the rotations of $\mathbb{S}^{2}$, then $T\circ R\circ T^{-1}$ runs over the maps of $\mathbb{C}\cup\{\infty\}$ of the form
$z\mapsto\frac{\alpha z+\beta}{-\overline{\beta}z+\overline{\alpha}},\quad(\alpha,\beta)\in\mathbb{C}^{2}\setminus\{(0,0)\}.$ (2.9)
The invariance by scaling allows to take $|\alpha|^{2}+|\beta|^{2}=1$. More information may be found in Needham (2023); Gel’fand et al. (1963). As a consequence, if $Z$ is the planar point process or Coulomb gas (2.1), then for every $(\alpha,\beta)$ as above,
$\frac{\alpha Z+\beta}{-\overline{\beta}Z+\overline{\alpha}}\overset{\mathrm{d}}{=}Z.$ (2.10)
It is convenient to think of $\mathbb{C}\cup\{\infty\}$ as the projective line $\mathbb{C}P^{1}=(\mathbb{C}^{2}\setminus\{(0,0)\})/\sim$, where the relation $\sim$ defining the quotient is given by $(z_{1},z_{2})\sim(w_{1},w_{2})$ if they are colinear : the equivalence class $[(z,w)]$ should be thought of as the complex line passing through the origin and $(z,w)$. A usual way to identify them is via the map $z\mapsto[(z,1)]$ which takes the metric $(1+|z|^{2})^{-2}|\mathrm{d}z|^{2}$ to the Fubini – Study metric [^5] We quickly recall a definition of the Fubini – Study metric. Fix $\ell=[(z,w)]$ thought of as a subspace of $\mathbb{C}^{2}$ and let us define a Hermitian metric on the tangent space $T_{\ell}\mathbb{C}P^{1}$. Noticing that $T_{\ell}\mathbb{C}P^{1}$ is canonically identified with the space of linear maps from $\ell$ to $\mathbb{C}^{2}/\ell$, the Hermitian metric on $\mathbb{C}^{2}$ induces one on $T_{\ell}\mathbb{C}P^{1}$. For another standard definition using a local potential see (Huybrechts, 2005, Example 3.1.9)., the metric of $\mathbb{C}P^{1}$ induced from the Hermitian product of $\mathbb{C}^{2}$. In this way, the invariance under rotations jumps out at us. Take any Hermitian product on $\mathbb{C}^{n\times n}$ and endow $(\mathbb{C}^{n\times n})^{2}$ with the product Hermitian product or, what is the same, endow $\mathbb{C}^{n\times n}\otimes\mathbb{C}^{2}$ with the tensor product of the Hermitian product of $\mathbb{C}^{n\times n}$ and the standard one of $\mathbb{C}^{2}$. Then, the map $(\mathbb{C}^{n\times n})^{2}\to(\mathbb{C}^{n\times n})^{2}$,
$(X,Y)\mapsto(\alpha X+\beta Y,-\overline{\beta}X+\overline{\alpha}Y),$
preserves the Hermitian product of $(\mathbb{C}^{n\times n})^{2}$ so that if $(X,Y)$ follows a standard Gaussian law in $(\mathbb{C}^{n\times n})^{2}$, its image $(\alpha X+\beta Y,-\overline{\beta}X+\overline{\alpha}Y)$ also does. This implies that $XY^{-1}\overset{\mathrm{d}}{=}(\alpha X+\beta Y)(-\overline{\beta}X+\overline{\alpha}Y)^{-1}=(\alpha XY^{-1}+\beta)(-\overline{\beta}XY^{-1}+\overline{\alpha})^{-1}$.
2.2. Kostlan observation and spectral radius
Following Kostlan (1992); Hough et al. (2009), if $Z_{n}=(Z_{n,1},\ldots,Z_{n,n})$ is a random vector of $\mathbb{C}^{n}$ with density (2.1), then the determinantal structure and rotational invariance in (2.1) give
$\{|Z_{n,1}|,\ldots,|Z_{n,n}|\}\overset{\mathrm{d}}{=}\{\xi_{n,1},\ldots,\xi_{n,n}\}$ (2.11)
where $\xi_{n,1},\ldots,\xi_{n,n}$ are independent and $\xi_{n,k}$ has density proportional to
$x\geq 0\mapsto x^{2k-1}\mathrm{e}^{-(n+1)Q(x)}=\frac{x^{2k-1}}{(1+x^{2})^{n+1}}.$ (2.12)
It follows that the random variable $\xi_{n,k}^{2}$ has density proportional to
$x\geq 0\mapsto\frac{x^{k-1}}{(1+x)^{n+1}}.$ (2.13)
We recognize a Beta prime (or inverse Beta or Beta of the second kind) distribution of density
$x\geq 0\mapsto\frac{\Gamma(a+b)}{\Gamma(a)\Gamma(b)}\frac{x^{a-1}}{(1+x)^{a+b}},$ (2.14)
with $a=k$ and b=nk+1b=n-k+1, which is also the law of B/(1B)B/(1-B) when $B\sim\mathrm{Beta}(a,b)$, and also the law of $G_{a}/G_{b}$ were $G_{a}\sim\mathrm{Gamma}(a,\lambda)$ and $G_{b}\sim\mathrm{Gamma}(b,\lambda)$ are independent and $\lambda>0$ is an arbitrary scale parameter.
Moreover, we also note at this step that $n/\xi_{n,n-k}^{2}\overset{\mathrm{d}}{\to}\mathrm{Gamma}(k,1)$ as $n\to\infty$. Following Chafaï and Péché (2014); Jiang and Qi (2017), it is possible to get from (2.11) the asymptotic fluctuations of the spectral radius, namely
$\frac{1}{\sqrt{n}}\rho(M^{\mathrm{Gin}})=\frac{1}{\sqrt{n}}\max_{1\leq k\leq n}|Z_{n,k}|\overset{\mathrm{d}}{=}\frac{1}{\sqrt{n}}\max_{1\leq k\leq n}\xi_{n,k}\xrightarrow[n\to\infty]{\mathrm{d}}\mathrm{Law}\Bigr(\max_{k\geq 1}\frac{1}{\sqrt{\gamma_{k}}}\Bigr)$ (2.15)
where ${(\gamma_{k})}_{k\geq 1}$ are independent with $\gamma_{k}\sim\mathrm{Gamma}(k,1)$. This is an alternative approach to the one using (3.4).
2.3. Equilibrium measure
Regarding high dimensional asymptotic analysis, almost surely,
$\mu_{M^{\mathrm{Gin}}}\overset{\mathrm{d}}{=}\frac{1}{n}\sum_{k=1}^{n}\delta_{Z_{n,k}}\xrightarrow[n\to\infty]{\text{weak}}\frac{\Delta Q(\left|\cdot\right|)}{2\pi}\mathrm{d}^{2}z=\frac{\mathrm{d}^{2}z}{\pi(1+|z|^{2})^{2}}=\nu,$ (2.16)
see for instance Hardy (2012). The average version of (2.16) is easy to check using the logarithmic potential and remains valid non asymptotically. Indeed, for all $z\in\mathbb{C}$, $B^{\mathrm{Gin}}$ and $A^{\mathrm{Gin}}-zB^{\mathrm{Gin}}$ are correlated, but by Gaussianity,
$(M^{\mathrm{Gin}}-z)B^{\mathrm{Gin}}=A^{\mathrm{Gin}}-zB^{\mathrm{Gin}}\overset{\mathrm{d}}{=}\sqrt{1+|z|^{2}}A^{\mathrm{Gin}}.$ (2.17)
Therefore, for all $z\in\mathbb{C}$, in $[-\infty,+\infty)$,
$\mathbb{E}\log|\det(M^{\mathrm{Gin}}-z)|+\mathbb{E}\log|\det(B^{\mathrm{Gin}})|=n\log\sqrt{1+|z|^{2}}+\mathbb{E}\log|\det(A^{\mathrm{Gin}})|.$ (2.18)
Finally, by applying the operator $\frac{1}{2\pi}\Delta$ in the sense of distributions, we obtain the stunning non-asymptotic formula
$\mathbb{E}\mu_{M^{\mathrm{Gin}}}=\frac{\Delta\log(1+|z|^{2})}{4\pi}\mathrm{d}^{2}z=\frac{\mathrm{d}^{2}z}{\pi(1+|z|^{2})^{2}}=\nu.$ (2.19)
Alternatively, we could use the fact that the uniform distribution on $\mathbb{S}^{2}$ is the unique distribution on $\mathbb{S}^{2}$ invariant by all rotations, together with the fact that its image by the stereographic projection $T$ is precisely $\nu$.
This formula is well known; see, for instance, Krishnapur (2009), Rogers (2010), and Forrester and Mays (2012). More generally, if $A$ and $B$ are Girko matrices, then $A-zB=\sqrt{1+|z|^{2}}C_{z}$ where $C_{z}$ is a Girko matrix, but with a law that depends on $z$ in general, except in the Gaussian case. Also the argument above works only in the Ginibre case.
3. Proof of Theorem 1.2 and Theorem 1.3
The key idea in a nutshell is as follows: if $A$ and $B$ are independent and have the same law, then
$M^{-1}\overset{\mathrm{d}}{=}M,\quad\text{which gives}\quad\rho_{\max}(M)\overset{\mathrm{d}}{=}\rho_{\max}(M^{-1})=\max_{\lambda\in\mathrm{spec}(M)}\frac{1}{|\lambda|}=\frac{1}{\min_{\lambda\in\mathrm{spec}(M)}|\lambda|}=\frac{1}{\mathrm{dist}(0,\mathrm{spec}(M))}.$ (3.1)
This remains valid when $A$ and $B$ do not have the same law, provided that in the right hand side of the equality in distribution symbols we modify $M$ by swapping the laws of $A$ and $B$, which is not a problem for our universality objective. In other words, the model $M$ is globally invariant in law by inversion. Geometrically, the inversion exchanges on the two-sphere the points $\infty$ and 0, the large scale behavior with the local scale behavior at the origin.
Proof of Theorem 1.2.
We will first study the case when $A$ and $B$ are Ginibre matrices so that $M=M^{\mathrm{Gin}}$. To simplify calculations we argue why it is enough to deal with the origin. Indeed, for any $\lambda_{0}\in\mathbb{C}$, we may take a rotation $R$ of $\mathbb{S}^{2}\simeq\mathbb{C}\cup\{\infty\}$ that maps 0 to $\lambda_{0}$. For instance, we could choose the explicit rotation
$R(z)=\frac{z+\lambda_{0}}{-\overline{\lambda_{0}}z+1}.$ (3.2)
For any such rotation $R$, we can show that $|\mathrm{det}(\mathrm{d}R_{0})|=(1+|\lambda_{0}|^{2})^{2}$ because it preserves the uniform measure on the sphere $(1+|z|^{2})^{-2}\mathrm{d}^{2}z$ but using the explicit $R$ from (3.2) we can directly calculate $R^{\prime}(0)=1+|\lambda_{0}|^{2}$ so that in this case $\mathrm{d}R_{0}=1+|\lambda_{0}|^{2}$. If we prefer to use a general rotation, using that $\mathrm{d}R_{0}$ is a multiplication by a complex number of squared norm $(1+|\lambda_{0}|^{2})^{2}$ we can also write $\mathrm{d}R_{0}=(1+|\lambda_{0}|^{2})e^{\mathrm{i}\theta}$ for some real $\theta$. Finally, by Lemma 5.3, if we consider $x_{0}=0$, $\varphi_{1}=R$ and $\varphi_{2}$ as the identity map, we get that
[…]
assuming the theorem holds for the origin. Since $\mathrm{Gin}_{\infty}$ is invariant under a multiplication by $e^{i\theta}$, we can forget the factor $e^{i\theta}$ so that $\mathrm{d}R_{0}(\mathrm{Gin}_{\infty})=(1+|\lambda_{0}|^{2})\mathrm{Gin}_{\infty}$. On the other hand, since $\{R(\lambda):\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}$ has the same law as $\{\lambda:\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}$, we get that
$\lim_{n\to\infty}\{\sqrt{n}(\lambda-\lambda_{0}):\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}=(1+|\lambda_{0}|^{2})\mathrm{Gin}_{\infty}.$
We suppose now that $\lambda_{0}=0$. Let $K_{n}$ be the kernel of the spectrum of $M^{\mathrm{Gin}}$ seen as a determinantal point process, given by (2.3). It follows that the spectrum of $\sqrt{n}M^{\mathrm{Gin}}$ forms a determinantal point process on $\mathbb{C}$ with kernel
$\frac{1}{n}K_{n}\Bigr(\frac{z}{\sqrt{n}},\frac{w}{\sqrt{n}}\Bigr)=\frac{\Bigr(1+\frac{z\overline{w}}{n}\Bigr)^{n-1}}{\pi\Bigr(1+\frac{|z|^{2}}{n}\Bigr)^{\frac{n+1}{2}}\Bigr(1+\frac{|w|^{2}}{n}\Bigr)^{\frac{n+1}{2}}}\xrightarrow[n\to\infty]{}\frac{\mathrm{e}^{z\overline{w}}}{\pi\mathrm{e}^{\frac{1}{2}|z|^{2}}\mathrm{e}^{\frac{1}{2}|w|^{2}}}=\sqrt{\gamma(z)\gamma(w)}\mathrm{e}^{z\overline{w}}\quad\text{where}\quad\gamma(z)=\frac{\mathrm{e}^{-|z|^{2}}}{\pi}.$ (3.3)
We recognize on the right-hand side the kernel of the infinite Ginibre $\mathrm{Gin}_{\infty}$ determinantal point process. Since the convergence is uniform with respect to $z$ and $w$ on compact subsets of $\mathbb{C}^{2}$, it follows that the determinantal point processes converge in distribution by, for instance, (Shirai and Takahashi, 2003, Proposition 3.10). In particular,
$\lim_{n\to\infty}\mathbb{E}\Bigr[F\Bigr(\int f(\sqrt{n}(z-z_{0}))\mathrm{d}\mu_{M^{\mathrm{Gin}}}(z)\Bigr)\Bigr]=\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\mu_{(1+|z_{0}|^{2})\mathrm{Gin_{\infty}}}\Bigr)\Bigr]$
for every bounded continuous function $F:\mathbb{R}\to\mathbb{R}$ and any compactly supported continuous function $f:\mathbb{C}\to\mathbb{R}$.
Let us now consider general $A$, $B$ and $M=AB^{-1}$ as in the hypothesis of Theorem 1.2. The replacement principle for local universality at any point, Theorem 4.1, tells us that we also have
$\lim_{n\to\infty}\mathbb{E}\Bigr[F\Bigr(\int f(\sqrt{n}(z-z_{0}))\mathrm{d}\mu_{M}(z)\Bigr)\Bigr]=\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\mu_{(1+|z_{0}|^{2})\mathrm{Gin_{\infty}}}\Bigr)\Bigr]$
but, for now, only for $F$ and $f$ smooth compactly supported functions. We can now apply Lemmas 5.1 and 5.2 from the last section to the sequence of point processes $(\mathcal{X}_{n})_{n\geq 1}$ given by $\mathcal{X}_{n}=\{\sqrt{n}(\lambda-z_{0}):\lambda\in\mathrm{spec}(M)\}$ and to the point process $\mathcal{X}$ being the scaled Ginibre process $(1+|z_{0}|^{2})\mathrm{Gin}_{\infty}$.
Proof of Theorem 1.3.
Let us first show the convergence of $\sqrt{n}\rho_{\mathrm{min}}(M)$. Since the map that takes a point configuration on $\mathbb{C}$ to the point configuration made of their norms is continuous and the map taking a point configuration on $0,\infty)$ to the minimum is also continuous, we get the convergence of the distance to the origin combining with Theorem 1.2. For a proof of the latter standard assertion, we may see (Butez and García-Zelada, 2022, Lemma A.1). Now an observation due to Kostlan, see Kostlan (1992) and (Hough et al., 2009, Th. 4.7.3 and Sec. 7.2), states that
$\{|\lambda|:\lambda\in\mathrm{Gin}_{\infty}\}\overset{\mathrm{d}}{=}\{\xi_{k}:k\geq 1\}$ (3.4)
where ${(\xi_{k})}_{k\geq 1}$ are independent random variables with $\xi_{k}^{2}\sim\mathrm{Gamma}(k,1)$ for all $k\geq 1$. If $\xi_{\min}=\min\{\xi_{k}:k\geq 1\}$,
$\sqrt{n}\rho_{\min}(M)\xrightarrow[n\to\infty]{\mathrm{law}}\xi_{\min}$
and we may calculate $\mathbb{P}(\xi_{\min}\geq r)$ as
$\mathbb{P}(\xi_{\min}\geq r)=\mathbb{P}\Bigr(\min_{k\geq 1}\xi_{k}\geq r\Bigr)=\prod_{k=1}^{\infty}\mathbb{P}(\xi_{k}\geq r)=\prod_{k=1}^{\infty}(1-F_{k})(r^{2})\quad\text{where}\quad F_{k}(x)=\mathbb{P}(\xi_{k}^{2}<x).$
Furthermore, the Gamma–Poisson duality gives $\mathbb{P}(\xi_{k}^{2}<x)=\mathbb{P}(\pi_{x}<k)=\mathrm{e}^{-x}\sum_{j=0}^{k-1}\frac{x^{j}}{j!}$ where $\pi_{x}\sim\mathrm{Poisson}(x)$.
Finally, for the convergence of $\rho_{\max}(M)/\sqrt{n}$, we notice that
$\rho_{\max}(AB^{-1})=\frac{1}{\rho_{\min}(BA^{-1})}$
so that it is a consequence of the convergence of the minimum for $BA^{-1}$. Notice the immediate but interesting fact that if $A$ and $B$ have the same law, the law of the matrix $M=AB^{-1}$ is invariant in law by inversion.
Alternatively, we could use the Kostlan observation on the spherical ensemble before passing to the limit, avoiding the usage of $\mathrm{Gin}_{\infty}$, see (2.15). ∎
4. Proof of Theorem 1.4
Theorem 1.4 is an immediate consequence of the following more general and technical result, used with the choice $f_{n}(z)=f(\sqrt{n}(z-z_{0}))$ for $f\in\mathcal{C}^{\infty}_{c}({\mathbb{C}}\to\mathbb{R})$ and $z_{0}\in{\mathbb{C}}$.
Theorem 4.1 (Comparison theorem).
Let $C_{0},C_{1}>0$ and $\nu>0$ be fixed and such that $\nu<10^{-4}c_{0}$ where $c_{0}$ is as in condition (C2).
Let $(f_{n})$} be a sequence of functions such that
$f_{n}\in\mathcal{C}_{c}^{2}({\mathbb{C}}\to\mathbb{R}),\qquad\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu},\qquad\mathrm{supp}(f_{n})\subset\{z\in\mathbb{C}:|z|\leq C_{0}\}.$ (4.1)
Let $F:\mathbb{R}\to\mathbb{R}$ be a function with uniformly bounded derivatives by $C_{1}$ up to fifth order, namely
$\max_{1\leq k\leq 5}\sup_{x\in{\mathbb{R}}}|F^{(k)}(x)|\leq C_{1}.$ (4.2)
Then, for all $0<\delta<c_{0}/400$, we have
$\Bigr|\mathbb{E}\Bigr[F\Bigr(\int f_{n}(z)\mathrm{d}\mu_{M}(z)\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f_{n}(z)\mathrm{d}\mu_{M^{\mathrm{Gin}}}(z)\Bigr)\Bigr]\Bigr|=O\big(n^{-\delta}\big).$
Theorem 4.1 proves universality in a broad sense, not only for gap distribution or spectral radius, but also for CLT with test functions satisfying (4.1). One can also use it to show the universality of convergence speed, as long as the optimal speed is much smaller than $n^{-\delta}$ for a sufficiently small $\delta>0$.
Before we present the proof of Theorem 4.1, we sketch an outline of it, in three steps:
  1. (1)
    Via the Hermitization trick introduced by Girko (1984), one reduces the spectral problem of $AB^{-1}$ near the point $z_{0}\in{\mathbb{C}}$ to studying the singular value statistics of AzBA-zB with $z$ near $z_{0}$, i.e.,
    $\mathcal{L}=\int f_{n}(x)\mathrm{d}\mu_{AB^{-1}}(x)=\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\log\det[(A-zB)(A-zB)^{*}]\mathrm{d}^{2}z,$ (4.3)
    see more details in (4.1)-(4.7) below. In particular, assuming $B$ is invertible, $z$ is an eigenvalue of $AB^{-1}$ is equivalent to that the smallest singular value of AzBA-zB is zero. This translates the eigenvalue problem of the non-Hermitian matrix into that of the Hermitian model $(A-zB)(A-zB)^{*}$ near zero.
  2. (2)
    To regularize the quantity on the right side of (4.3), we introduce a parameter $\eta_{0}$ slightly below the typical size of the smallest singular value of AzBA-zB near zero and show that (see Lemma 4.2)
    $\mathcal{L}\approx\mathcal{L}^{\eta_{0}}=$ $\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\log\det\big[(A-zB)(A-zB)^{*}+\eta_{0}^{2}\big]\mathrm{d}^{2}z$
    $\approx$ $-\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\int_{\eta_{0}}^{T}\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)\mathrm{d}\eta\mathrm{d}^{2}z,\qquad\qquad\eta_{0}=n^{-1-\epsilon},\quad T=n^{100},$ (4.4)
    for a small $\epsilon>0$, where the approximation holds in the sense of absolute mean, and $G^{z}$ is the resolvent (or Green function) of the Hermitised matrix of AzBA-zB as defined in (4.7))-(4.8) below. This step relies on the effective lower bound of the smallest singular value of Girko matrices from Jain et al. (2021) and the local law estimates for the Green function from Ajanki et al. (2014). It is worth noting that the polynomial lower bound for the smallest singular value obtained in Bordenave and Chafaï (2012) and inspired from Tao and Vu (2010) is not enough for our purposes regarding very tiny singular values, see (4.1).
  3. (3)
    From (2), it then suffices to show that under the fourth moment condition (C2) (see Lemma 4.3),
    $\mathbb{E}[F(\mathcal{L}^{\eta_{0}})]=\mathbb{E}^{\mathrm{Gin}}[F(\mathcal{L}^{\eta_{0}})]+o(1),$
    where $\mathbb{E}^{\mathrm{Gin}}$ indicates the corresponding expectation with $A$ and $B$ being Ginibre matrices. This step relies on the so-called long-time Green function comparison theorem (GFT). To achieve this, we let the matrix entries of $A$ and $B$ run independent Ornstein-Uhlenbeck dynamics. Then the corresponding matrix flow $A_{t}-zB_{t}$ interpolates between AzBA-zB and the Ginibre counterpart $A^{\mathrm{Gin}}-zB^{\mathrm{Gin}}$. We use Itô’s formula and cumulant expansion formula to compute the time derivative of $\mathbb{E}[F(\mathcal{L}_{t}^{\eta_{0}})]$ and then use the local law estimates for the Green function and the fourth moment condition (C2) to show the smallness of $\mathrm{d}\mathbb{E}[F(\mathcal{L}_{t}^{\eta_{0}})]$ along the interpolating flow.
4.1. Proof of Theorem 4.1
In this section, we normalize both Girko matrices $A$ and $B$ by $1/\sqrt{n}$, i.e., ${\tilde{A}}=A/\sqrt{n}$ and ${\tilde{B}}=B/\sqrt{n}$. It is easy to check that $M=AB^{-1}={\tilde{A}}{\tilde{B}}^{-1}$. We also set $M^{\mathrm{Gin}}={\tilde{A}}^{\mathrm{Gin}}({\tilde{B}}^{\mathrm{Gin}})^{-1}=A^{\mathrm{Gin}}(B^{\mathrm{Gin}})^{-1}$ with ${\tilde{A}}^{\mathrm{Gin}}=A^{\mathrm{Gin}}/\sqrt{n}$, ${\tilde{B}}^{\mathrm{Gin}}=B^{\mathrm{Gin}}/\sqrt{n}$ being normalized Ginibre matrices. For conventional consistency, we always write $\mathrm{Gin}$ as a superscript to denote the special Gaussian case.
Using that $\Delta_{z}\log|z|=2\pi\delta_{0}$ and applying integration by parts twice, for any $f\in\mathcal{C}^{2}_{c}({\mathbb{C}}\rightarrow{\mathbb{R}})$, we have
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=$ $\frac{1}{2\pi}\int\Delta_{z}f_{n}(z)\log|\det({\tilde{A}}{\tilde{B}}^{-1}-z)|$
$=$ $\frac{1}{2\pi}\int\Delta_{z}f_{n}(z)\big(\log|\det({\tilde{A}}-z{\tilde{B}})|-\log|\det({\tilde{B}})|\big)\mathrm{d}^{2}z,$ (4.5)
[…]
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=\frac{1}{2\pi}\int\Delta_{z}f_{n}(z)\log|\det(X^{z})|\mathrm{d}^{2}z,\qquad X^{z}={\tilde{A}}-z{\tilde{B}}\in{\mathbb{C}}^{n\times n}.$ (4.6)
By the Hermitization trick of Girko (1984), we write
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\log[\det H^{z}|\mathrm{d}^{2}z,\qquad H^{z}=\begin{pmatrix}0&X^{z}\\ (X^{z})^{*}&0\end{pmatrix}\in{\mathbb{C}}^{2n\times 2n}.$ (4.7)
The eigenvalues of the Hermitised matrix $H^{z}$ are then denoted by $\big(\lambda^{z}_{\pm i}\big)_{i=1}^{n}$ with $\lambda^{z}_{-i}=-\lambda^{z}_{i}$ and $0\leq\lambda^{z}_{i}\leq\lambda^{z}_{i+1}$ for $1\leq i\leq n$. Note that the non-negative eigenvalues $\big(\lambda^{z}_{i}\big)_{i=1}^{n}$ are also the singular values of $X^{z}=(A-zB)/\sqrt{n}$. We further define the resolvent (or Green function) of $H^{z}~(z\in\mathbb{C})$ by
$G^{z}(w)=(H^{z}-w)^{-1},\qquad\quad w\in\mathbb{C}\setminus\mathbb{R}.$ (4.8)
Using that $\partial_{\eta}\log(\lambda^{2}+\eta^{2})=\frac{2\eta}{\lambda^{2}+\eta^{2}}$, we obtain from (4.7) that
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=-\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\int_{0}^{T}\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)\mathrm{d}\eta\mathrm{d}^{2}z+\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\log|\det(H^{z}-\mathrm{i}T)|\mathrm{d}^{2}z,$ (4.9)
where we chose $T=n^{100}$. Note that
$\log|\det(H^{z}-\mathrm{i}T)|-2n\log T=\sum_{j}\log\left(1+\Big(\frac{\lambda_{j}^{z}}{T}\Big)^{2}\right)\leq\frac{\mathrm{Tr}(H^{z})^{2}}{T^{2}}\prec n^{-199},$ (4.10)
where we used that $\log(1+|x|)\leq|x|$ and, for the stochastic domination, we used (C3) to get that
$\mathbb{P}(\mathrm{Tr}(X^{z})(X^{z})^{*}\geq n^{\xi}n)=\mathbb{P}\Bigr(\sum_{i,j=1}^{n}|a_{ij}-zb_{ij}|^{2}\geq n^{\xi}n^{2}\Bigr)\leq\sum_{i,j=1}^{n}\mathbb{P}(|a_{ij}-zb_{ij}|^{2}\geq n^{\xi})\leq 2^{k-1}nn^{-k\xi}D_{k}$ (4.11)
which implies $\mathrm{Tr}(X^{z})(X^{z})^{*}\prec n$. Recall that $f_{n}\in\mathcal{C}^{2}_{c}({\mathbb{C}})$ satisfies $\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu}$ with $\nu<10^{-4}c_{0}$. Hence we have
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=-\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\int_{0}^{T}\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)\mathrm{d}\eta\mathrm{d}^{2}z+O_{\prec}(n^{-100}).$ (4.12)
We next split the above $\eta$-integral into two parts, i.e.,
$I^{\leq\eta_{0}}+I^{\geq\eta_{0}}=-\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\Big(\int_{0}^{\eta_{0}}+\int_{\eta_{0}}^{T}\Big)\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)\mathrm{d}\eta\mathrm{d}^{2}z,\quad\qquad\eta_{0}=n^{-1-\epsilon},$ (4.13)
for some small $\epsilon>0$ to be fixed later. The lemma below shows that the first $\eta$-integral $I^{\leq\eta_{0}}$ is negligible in the sense of absolute mean.
Lemma 4.2.
Fix any small $\epsilon>0$ and set $\eta_{0}=n^{-1-\epsilon}$. Then the following holds
$\mathbb{E}\big|I^{\leq\eta_{0}}\big|=O\big(n^{-\epsilon/2+\nu}\big).$
Recall that $c_{0}>0$ is a small constant in (C2) and $\nu<10^{-4}c_{0}$ from (4.1). Hence we can choose a small $\epsilon>0$ such that $100\nu<\epsilon<c_{0}/100$, where the upper restriction of $\epsilon$ will be used in Lemma 4.3 below. Note that $F$ has a bounded derivative as in (4.2). Then by a simple Taylor expansion and using Lemma 4.2 with $\nu<\epsilon/100$, we have
$\mathbb{E}\Big[F\Big(\int f_{n}(x)\mathrm{d}\mu_{M}(x)\Big)\Big]=\mathbb{E}\big[F(I^{\geq\eta_{0}})\big]+O\big(n^{-\epsilon/4}\big).$ (4.14)
Since the Ginibre matrix is also a Girko matrix, the above also applies to $M^{\mathrm{Gin}}=A^{\mathrm{Gin}}(B^{\mathrm{Gin}})^{-1}$ with $\tilde{A}$} and $\tilde{B}$} replaced by $A^{\mathrm{Gin}}$ and $B^{\mathrm{Gin}}$. It then suffices to compare the expectation $\mathbb{E}\big[F(I^{\geq\eta_{0}})\big]$ with its Ginibre counterpart $\mathbb{E}^{\mathrm{Gin}}\big[F(I^{\geq\eta_{0}})\big]$.
Lemma 4.3.
Recall that $\nu<10^{-4}c_{0}$ from (4.1). Fix any $0<\epsilon<c_{0}/100$ and let $\eta_{0}=n^{-1+\epsilon}$. Then
$\Big|\mathbb{E}\big[F(I^{\geq\eta_{0}})\big]-\mathbb{E}^{\mathrm{Gin}}\big[F(I^{\geq\eta_{0}})\big]\Big|=O(n^{-c_{0}/4}),$ (4.15)
where we used $\mathbb{E}^{\mathrm{Gin}}$ to denote the corresponding expectation with $\tilde{A}$} and $\tilde{B}$} being Ginibre matrices.
Therefore, Theorem 4.1 follows directly from (4.14) for both matrices $M$ and $M^{\mathrm{Gin}}$ in combination with Lemma 4.3 for $\epsilon<c_{0}/100$, where $c_{0}>0$ is small constant from (C2).
The remaining part of this section is to prove Lemma 4.2 and Lemma 4.3.
Proof of Lemma 4.2.
Recall that $H^{z}$ in (4.7) is the Hermitised matrix of $X^{z}=\widetilde{A}-z\widetilde{B}$, and $G^{z}$ is its resolvent defined in (4.8). For each fixed $z\in{\mathbb{C}}$, $X^{z}=(x_{ij})$ is still a Girko matrix with $\mathbb{E}[x_{ij}]=0$, $\mathbb{E}[(x_{ij})^{2}]=0$ and $\mathbb{E}[|x_{ij}|^{2}]=(1+|z|^{2})/n$. In particular if we set $z=0$, then $X^{z=0}$ becomes a standard Girko matrix with variance $n^{-1}$. From (Ajanki et al., 2014, Theorem 1.1) (see also Cacciapuoti et al. (2013); Bourgade et al. (2014)), fixing any small $\xi>0$, the following local laws for the resolvent $G^{z}$ with $z=0$
$\Big|\big\langle\mathbf{x},\big(G^{z=0}(w)-m_{sc}(w)\mathbf{y}\big\rangle\Big|\prec\sqrt{\frac{1}{n\eta}},\qquad\Big|\frac{1}{n}\mathrm{Tr}G^{z=0}(w)-m_{sc}(w)\Big|\prec\frac{1}{n\eta},$ (4.16)
hold uniformly for any $|w|\leq 10$ and $\eta=|\operatorname{Im}w|\geq n^{-1+\xi}$ and for any deterministic unit vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$, where $m_{sc}$ is the Cauchy – Stieltjes transform of the semicircle distribution $\rho_{sc}(x)=\frac{1}{2\pi}\sqrt{(4-x^{2})_{+}}$. In general for any fixed $|z|\leq C_{0}$, by a simple rescaling, the following local laws for $G^{z}$
$\Big|\big\langle\mathbf{x},\big(G^{z}(w)-m^{z}(w)\mathbf{y}\big\rangle\Big|\prec\sqrt{\frac{1}{n\eta}},\qquad\Big|\frac{1}{n}\mathrm{Tr}G^{z}(w)-m^{z}(w)\Big|\prec\frac{1}{n\eta},$ (4.17)
hold uniformly for any $|w|\leq 10$ and $\eta=|\operatorname{Im}w|\geq n^{-1+\xi}$, with the deterministic function $m^{z}$ given by
$m^{z}(w)=\frac{1}{\sqrt{1+|z|^{2}}}m_{sc}\Big(\frac{w}{\sqrt{1+|z|^{2}}}\Big),\qquad w\in{\mathbb{C}}\setminus{\mathbb{R}}.$ (4.18)
It is easy to check that $m^{z}$ is uniformly bounded by some constant for any $|w|\leq 10$ and $|z|\leq C_{0}$. As a direct consequence of (4.17), we obtain the following estimate on the eigenvalue counting function $N(E_{1},E_{2})=\big\{\lambda_{j}\in\mathrm{Spec}(H^{z}):E_{1}\leq\lambda_{j}\leq E_{2}\big\}$:
$\big|N(E_{1},E_{2})-N\int_{E_{1}}^{E_{2}}\rho^{z}(x)\mathrm{d}x\big|\prec 1\qquad E_{1}<E_{2},$ (4.19)
where $\rho^{z}$ is the probability distribution whose Cauchy – Stieltjes transform is given by (4.18). The proof of (4.19) is quite standard (see similar arguments used to prove (Benaych-Georges and Knowles, 2017, Theorem 2.8)), so we omit it. We remark that though the above bound holds with very high probability for each fixed $z$, they can be extended to very high probability estimates simultaneously for all $|z|\leq C_{0}$ by a grid argument and union bound, since $G^{z}$ is Lipschitz continuous in $z$ and $\|G^{z}(\mathrm{i}\eta)\|\leq\eta^{-1}$. Without specific mention, all the estimates below hold simultaneously for all $|z|\leq C_{0}$.
Now we are ready to prove Lemma 4.2. By a simple spectral decomposition of $H^{z}$, we have
$I^{\leq\eta_{0}}=-\frac{1}{2\pi}\int\Delta_{z}f_{n}(z)\Big[\int_{0}^{\eta_{0}}\sum_{j=1}^{n}\frac{\eta}{(\lambda_{j}^{z})^{2}+\eta^{2}}\mathrm{d}\eta\Big]\mathrm{d}^{2}z,\qquad\eta_{0}=n^{-1-\epsilon},$ (4.20)
where $\big(\lambda^{z}_{j}\big)_{j=1}^{n}$ are the ordered non-negative eigenvalues of $H^{z}$, that are also singular values of $X^{z}$. Let $L>0$ be a large number to be fixed later. Set $\eta_{L}=n^{-L}$ and $\eta_{2}=n^{-1-\epsilon/2}$. Taking the expectation, we split the above sum of eigenvalues into three parts as below:
$\mathbb{E}\Big[\int_{0}^{\eta_{0}}\sum_{j=1}^{n}\frac{\eta}{(\lambda^{z}_{j})^{2}+\eta^{2}}\mathrm{d}\eta\mathrm{d}^{2}z\Big]=$ $\mathbb{E}\Big[\int_{0}^{\eta_{0}}\sum_{\lambda^{z}_{j}\geq\eta_{2}}\frac{\eta}{(\lambda_{j}^{z})^{2}+\eta^{2}}\mathrm{d}\eta\mathrm{d}^{2}z\Big]$
$+\mathbb{E}\Big[\Big(\sum_{\eta_{L}\leq\lambda^{z}_{j}\leq\eta_{2}}+\sum_{\lambda^{z}_{j}\leq\eta_{L}}\Big)\log\Big(1+\frac{\eta_{0}^{2}}{(\lambda^{z}_{j})^{2}}\Big)\Big],$ (4.21)
where we also used that $\partial_{\eta}\log(\lambda^{2}+\eta^{2})=\frac{2\eta}{\lambda^{2}+\eta^{2}}$ in the last line. For the first line of (4.1), we have
$\mathbb{E}\Big[\int_{0}^{\eta_{0}}\sum_{\lambda_{j}\geq\eta_{2}}\frac{\eta}{\lambda_{j}^{2}+\eta^{2}}\mathrm{d}\eta\Big]\leq$ $2\mathbb{E}\Big[\int_{0}^{\eta_{0}}\sum_{\lambda_{j}\geq\eta_{2}}\frac{\eta}{\lambda_{j}^{2}+\eta_{2}^{2}}\mathrm{d}\eta\Big]$
$\leq$ $\mathbb{E}\Big[\sum_{j=1}^{n}\frac{\eta_{0}^{2}}{(\lambda^{z}_{j})^{2}+\eta_{2}^{2}}\Big]=\frac{\eta_{0}^{2}}{2\eta_{2}}\mathbb{E}\big[\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta_{2})\big].$ (4.22)
Note that the function $\eta\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)$ is non-decreasing in $\eta>0$. Choosing $\eta_{3}=n^{-1+\xi}$ for a small $\xi<\epsilon/100$,
$\mathbb{E}\Big[\int_{0}^{\eta_{0}}\sum_{\lambda_{j}\geq\eta_{2}}\frac{\eta}{\lambda_{j}^{2}+\eta^{2}}\mathrm{d}\eta\Big]\leq\frac{\eta_{0}^{2}}{\eta_{2}}\frac{\eta_{3}}{\eta_{2}}\mathbb{E}\big[\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta_{3})\big]=O(n^{-\epsilon/2}),$ (4.23)
where we used the local law in (4.17) and (4.18) for $\eta_{3}=n^{-1+\xi}$.
We next estimate the last line of (4.1). Note that the non-negative eigenvalues $\big(\lambda^{z}_{i}\big)_{i=1}^{n}$ of $H^{z}$ coincide with the singular values of $X^{z}=(A-zB)/\sqrt{n}$, where $A$ and $B$ are independent Girko matrices satisfying (C1)-(C3). Recall from (Jain et al., 2021, Lemma 3.4) (see also Tikhomirov (2020); Erdős and Ji (2024)) that there exists a constant $C$ depending on $C_{0}$ and $D_{0}$ from (C1) such that the smallest singular value $\lambda_{1}^{z}$ of $(A-zB)/\sqrt{n}$ with $|z|\leq C_{0}$ satisfies
$\mathbb{P}\big(\lambda^{z}_{1}\leq t\big)\leq Cn^{2}t^{2},\qquad\forall t>0,\quad|z|\leq C_{0}.$ (4.24)
Moreover, from the bound (4.19) on the eigenvalue counting function near the origin, we obtain that
$\#\big\{j\in[n]:\lambda^{z}_{j}\leq n^{-1}\big\}\leq n^{\xi^{\prime}},$ (4.25)
for any small $\xi^{\prime}>0$, with very high probability. Using (4.24) for $t=n^{-\epsilon/2}$ and (4.25) for $\xi^{\prime}<\epsilon/100$, the first sum in the last line of (4.1) is bounded by
$\mathbb{E}\Big[\sum_{\eta_{L}\leq\lambda^{z}_{j}\leq\eta_{2}}\log\Big(1+\frac{\eta_{0}^{2}}{(\lambda^{z}_{j})^{2}}\Big)\Big]\leq n^{\xi^{\prime}}(\log n)\mathbb{P}\big(\lambda^{z}_{1}\leq\eta_{2}\big)=O(n^{-\epsilon/2}).$ (4.26)
For the last part in (4.1) with tiny eigenvalues $\lambda^{z}_{j}\leq\eta_{L}=n^{-L}$, using that $\log(1+|x|)\leq 2\log|x|$ for $|x|>10$ and the tail bound for the smallest singular value in (4.24), we obtain
$\mathbb{E}\Big[\sum_{\lambda^{z}_{j}\leq\eta_{L}}\log\Big(1+\frac{\eta_{0}^{2}}{(\lambda^{z}_{j})^{2}}\Big)\Big]\leq$ $4\mathbb{E}\Big[\sum_{\lambda^{z}_{j}\leq\eta_{L}}-\log(\lambda^{z}_{j})\Big]\leq 4n\mathbb{E}\Big[-\log\big(\lambda^{z}_{1}\big)\mathbf{1}_{\lambda^{z}_{1}\leq\eta_{L}}\Big]$
$=$ $4n\int_{L\log n}^{\infty}\mathbb{P}\big(\lambda^{z}_{1}\leq\mathrm{e}^{-s}\big)\mathrm{d}s=O(n^{-100}),$ (4.27)
by choosing a sufficiently large $L>0$ depending on the constants $C_{0}$ and $D_{0}>0$ from (C1).
Summing up (4.23), (4.26) and (4.1), we conclude from (4.20) and (4.1) that
$\mathbb{E}|I^{\leq\eta_{0}}|\leq\frac{1}{2\pi}\int\big|\Delta_{z}f_{n}(z)\big|\mathbb{E}\Big[\int_{0}^{\eta_{0}}\sum_{j=1}^{n}\frac{\eta}{(\lambda_{j}^{z})^{2}+\eta^{2}}\mathrm{d}\eta\Big]\mathrm{d}^{2}z=O\big(n^{-\epsilon/2+\nu}\big),$ (4.28)
where we also used that $\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu}$. This finished the proof of Lemma 4.2.
Proof of Lemma 4.3.
Define the following the Ornstein – Uhlenbeck matrix flow
$\mathrm{d}{\tilde{A}}_{t}=-\frac{1}{2}{\tilde{A}}_{t}\mathrm{d}t+\frac{\mathrm{d}\mathcal{M}^{(1)}_{t}}{\sqrt{n}},\qquad\mathrm{d}{\tilde{B}}_{t}=-\frac{1}{2}{\tilde{B}}_{t}\mathrm{d}t+\frac{\mathrm{d}\mathcal{M}^{(2)}_{t}}{\sqrt{n}},$ (4.29)
with initial ensembles ${\tilde{A}}_{0}={\tilde{A}}=A/\sqrt{n}$ and ${\tilde{B}}_{0}={\tilde{B}}=B/\sqrt{n}$, where $A,B$ are independent Girko matrices satisfying (C1)(C3), and $\mathcal{M}^{(1)}_{t}$ and $\mathcal{M}^{(2)}_{t}$ are two independent $n\times n$ matrices whose entries are i.i.d. standard complex Brownian motions. Then for any $z\in{\mathbb{C}}$ we define $X^{z}_{t}={\tilde{A}}_{t}-z{\tilde{B}}_{t}$ as in (4.6). In distribution, it satisfies
$\mathrm{d}X^{z}_{t}=-\frac{1}{2}X^{z}_{t}\mathrm{d}t+\sqrt{\frac{1+|z|^{2}}{n}}\mathrm{d}\mathcal{M}_{t},\qquad X^{z}_{0}=\frac{A-zB}{\sqrt{n}},$ (4.30)
where $\mathcal{M}_{t}$ is a standard complex Brownian motion matrix. For notational simplicity, we may often drop the dependence on $z$ or $t$ and write $X_{t}^{z}=(x_{ij})_{i,j\in[2n]}$. We use $X^{z}_{t}$ to construct the Hermitised matrix $H^{z}_{t}=(h_{ij})_{i,j\in[2n]}$ as in (4.7) and define the corresponding Green function $G^{z}_{t}(w)=(H^{z}_{t}-w)^{-1}$ as in (4.8). In particular,
$h_{aB}=\overline{h_{Ba}}=x_{a,B-n},\quad\qquad h_{Ba}=\overline{h_{aB}},\quad\qquad a\in[n],~B\in[n,2n],$ (4.31)
and all the other components of $H^{z}_{t}$ are equal to zero. Note that $(x_{ij})_{i,j\in[n]}$ are time-dependent random variables which also depend on $z$. Recall that the $(p,q)$-cumulant of a complex-valued random variable $\chi$ with finite moments are defined to be
$\kappa^{(p,q)}(\chi)=(-\mathrm{i})^{p+q}\Big(\frac{\partial^{p+q}}{\partial s^{p}\partial t^{q}}\log\mathbb{E}\big[\mathrm{e}^{\mathrm{i}s\chi+\mathrm{i}t\overline{\chi}}\big]\Big)\Big|_{s,t=0}.$ (4.32)
From (C2)-(C3), it is easy to check that second cumulants of $(x_{ij})_{i,j\in[n]}$ are invariant with time
$\kappa^{(1,1)}(x_{ij})=\frac{1+|z|^{2}}{n},\qquad\kappa^{(0,2)}(x_{ij})=\kappa^{(2,0)}(x_{ij})=0,$ (4.33)
and the higher-order cumulants are bounded by
$|\kappa^{(p,q)}(x_{ij})|\lesssim n^{-2-c_{0}},\qquad p+q=3,4,\qquad|\kappa^{(p,q)}(x_{ij})|\lesssim n^{-\frac{p+q}{2}},\qquad p+q\geq 5,$ (4.34)
uniformly for any $t\geq 0$ and $|z|\leq C_{0}$.
Recall the definition of $I^{\geq\eta_{0}}$ in (4.13). Note that $\mathrm{Tr}G^{z}(\mathrm{i}\eta)$ is purely imaginary due to the block structure of $H^{z}$ in (4.7), so we have $\mathrm{Tr}G^{z}(\mathrm{i}\eta)=\mathrm{i}\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)$. Then we define the following short-hand
$I^{\geq\eta_{0}}_{t}=\frac{\mathrm{i}}{4\pi}\int\Delta_{z}f_{n}(z)\int_{\eta_{0}}^{T}\mathrm{Tr}G^{z}_{t}(\mathrm{i}\eta)\mathrm{d}\eta\mathrm{d}^{2}z,$ (4.35)
with $T=n^{100}$ and $\eta_{0}=n^{-1-\epsilon}$ for $0<\epsilon<c_{0}/100$.
We then use Itô’s formula to derive the dynamics of $F(I^{\geq\eta_{0}}_{t})$ as a function of $X^{z}_{t}=\{x_{ij}\}_{i,j\in[n]}$ given by (4.30). By direct computations and the relation (4.31), we obtain
$\frac{\mathrm{d}F(I^{\geq\eta_{0}}_{t})}{\mathrm{d}t}=\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\left(-\frac{1}{2}h_{aB}\frac{\partial F(I^{\geq\eta_{0}}_{t})}{\partial h_{aB}}-\frac{1}{2}\overline{h_{aB}}\frac{\partial F(I^{\geq\eta_{0}}_{t})}{\partial\overline{h_{aB}}}+\frac{1+|z|^{2}}{n}\frac{\partial^{2}F(I^{\geq\eta_{0}}_{t})}{\partial h_{aB}\partial\overline{h_{aB}}}\right)\mathrm{d}t+\mathrm{d}M,$ (4.36)
where $\mathrm{d}M$ denotes the corresponding diffusion term with vanishing expectation, due to the martingale property of Brownian motions. We next take the expectation of the remaining drift terms in (4.36). Note that the differentiation rules below are obtained from the definition of resolvent in (4.8):
$\frac{\partial(G^{z}_{t})_{xy}}{\partial h_{aB}}=-(G^{z}_{t})_{xa}(G^{z}_{t})_{By},\qquad\frac{\partial(G^{z}_{t})_{xy}}{\partial\overline{h_{aB}}}=-(G^{z}_{t})_{xB}(G^{z}_{t})_{ay},$ (4.37)
for any $x,y\in[2n]$, $a\in[n]$, and $B\in[n,2n]$. Recall that $\kappa^{(p+1,q)}(h_{aB})$ and $\kappa^{(p+1,q)}(\overline{h_{aB}})$ are the $(p+1,q)$-cumulants of the complex valued random variables from (4.31). Applying the cumulant expansion formula (for a formal statement, see (He and Knowles, 2017, Lemma 7.1)) to the expectation of the first two terms in (4.36), the resulting second-order cumulant terms from (4.33) will cancel precisely with the last drift term in (4.36). This key cancellation is a consequence of the fact that the second moments of the matrix entries of $X^{z}_{t}$ do not change with time. Hence, we start with the third-order terms ($p+q+1=3$) in the cumulant expansions, i.e.,
$\frac{\mathrm{d}\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t})\big]}{\mathrm{d}t}=$ $-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\left(\sum_{p+q+1=3}^{K_{0}}\frac{\kappa^{(p+1,q)}(h_{aB})}{p!q!}\mathbb{E}\left[\frac{\partial^{p+q+1}F(I^{\geq\eta_{0}}_{t})}{\partial h_{aB}^{p+1}\partial\overline{h_{aB}}^{q}}\right]\right)$
$-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\left(\sum_{p+q+1=3}^{K_{0}}\frac{\kappa^{(p+1,q)}(\overline{h_{aB}})}{p!q!}\mathbb{E}\left[\frac{\partial^{p+q+1}F(I^{\geq\eta_{0}}_{t})}{\partial\overline{h_{aB}}^{p+1}\partial h_{aB}^{q}}\right]\right)+O(n^{-10}),$ (4.38)
where we stopped the expansion at a sufficiently high order with $p+q+1=K_{0}$ such that the truncation error is bounded by $O(n^{-10})$, using the cumulant bounds in (4.34), the differentiation rule in (4.37), the deterministic bound $\|G_{t}^{z}(\mathrm{i}\eta)\|\leq\eta^{-1}$, together with the local law estimates in (4.48) below. Such truncation argument is standard and frequently used in previous works e.g., He and Knowles (2017); Schnelli and Xu (2022); Cipolloni et al. (2024b), so we omit the details. We remark that the last error term holds uniformly for any $|z|\leq C_{0}$ and $0\leq t\leq 100\log n$.
To estimate the third and higher order terms with $3\leq p+q+1\leq K_{0}$, we introduce the following lemma, whose proof is postponed till the end of this section.
Lemma 4.4.
Recall $I^{\geq\eta_{0}}_{t}$ in (4.35) with $\eta_{0}=n^{-1-\epsilon}$ and compactly supported $f_{n}$ satisfying $\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu}$. Then for any $k=k_{1}+k_{2}\geq 1$, we have
$\sup_{t\in[0,t_{0}]}\Big|\frac{\partial^{k}I^{\geq\eta_{0}}_{t}}{\partial(h_{aB})^{k_{1}}(\overline{h_{aB}})^{k_{2}}}\Big|\prec n^{2k\epsilon+\nu},\qquad t_{0}=100\log n.$ (4.39)
Using the cumulant bounds in (4.34) and the differentiation bounds in (4.39) to estimate the third and higher order terms on the right side of (4.1), we obtain that
$\sup_{t\in[0,t_{0}]}\Big|\frac{\mathrm{d}\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t})\big]}{\mathrm{d}t}\Big|=O(n^{-c_{0}/2}),$ (4.40)
where we also used that $\epsilon<c_{0}/100$ and $\nu<10^{-4}c_{0}$. Integrating (4.40) over $t\in[0,t_{0}]$ with $t_{0}=100\log n$ we obtain
$\Big|\mathbb{E}\big[F(I^{\geq\eta_{0}}_{0})\big]-\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t_{0}})\big]\Big|=O(n^{-c_{0}/4}).$ (4.41)
Note that $H^{z}_{t}$ is defined as in (4.7) with the time dependent matrix $X^{z}_{t}$ from (4.30), which in distribution equals to
$X^{z}_{t}\stackrel{{\scriptstyle{\rm d}}}{{=}}\mathrm{e}^{-\frac{t}{2}}X^{z}+\sqrt{1-\mathrm{e}^{-t}}X^{\mathrm{Gin}},\qquad t\geq 0,$
with $X^{z}_{0}=X^{z}=(A-zB)/\sqrt{n}$ and $X^{z}_{\infty}=X^{\mathrm{Gin}}$ being the Gaussian counterpart which is independent of $X^{z}$. Using the resolvent identity, we obtain
$\|G^{z}_{t_{0}}(\mathrm{i}\eta)-G^{z}_{\infty}(\mathrm{i}\eta)\|\leq\|G^{z}_{t_{0}}(\mathrm{i}\eta)\|\|G^{z}_{\infty}(\mathrm{i}\eta)\|\|X^{z}_{t_{0}}-X^{z}_{\infty}\|\prec n^{-50},\qquad\eta\geq\eta_{0},$ (4.42)
where we also used that $\|G(\mathrm{i}\eta)\|\leq\eta^{-1}$ and $|x_{ij}|\prec n^{-1/2}$ from condition (C3). Thus using that $F$ has uniformly bounded derivative and $\int|\Delta_{z}f(z)|\mathrm{d}^{2}z\leq n^{\nu}$ with $\nu<10^{-4}c_{0}$, we obtain from (4.35) that
$\Big|\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t_{0}})\big]-\mathbb{E}\big[F(I^{\geq\eta_{0}}_{\infty})\big]\Big|=O(n^{-10}).$ (4.43)
Combining (4.41) with (4.43), we conclude
$\Big|\mathbb{E}\big[F(I^{\geq\eta_{0}})\big]-\mathbb{E}^{\mathrm{Gin}}\big[F(I^{\geq\eta_{0}})\big]\Big|=\Big|\mathbb{E}\big[F(I^{\geq\eta_{0}}_{0})\big]-\mathbb{E}\big[F(I^{\geq\eta_{0}}_{\infty})\big]\Big|=O(n^{-c_{0}/4}).$ (4.44)
We hence finish the proof of Lemma 4.3. ∎
We end this section with the proofs of Lemma 4.4.
Proof of Lemma 4.4.
Recall $I^{\geq\eta_{0}}_{t}$ in (4.35). Using (4.37) and $G^{2}=\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}\eta}G(\mathrm{i}\eta)$, we further have
$\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h_{aB}}=$ $-\frac{\mathrm{i}}{4\pi}\int\Delta_{z}f_{n}(z)\int_{\eta_{0}}^{T}\big[\big(G_{t}^{z}(\mathrm{i}\eta)\big)^{2}\big]_{Ba}\mathrm{d}\eta\mathrm{d}^{2}z=\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{Ba}\mathrm{d}^{2}z+O_{\prec}(n^{-10}),$ (4.45)
where we also used that $\|G(\mathrm{i}\eta)\|\leq\eta^{-1}$ for $\eta=T=n^{100}$. Note that the variances of entries of $X^{z}_{t}$ do not change with time (see (4.33)). Hence, for each fixed $t\geq 0$ and $|z|\leq C_{0}$, the local law as in (4.17) also holds true for $G^{z}_{t}$ with very high probability. To estimate the right side of (4.45) with $\eta_{0}=n^{-1-\epsilon}$, we next extend the estimates in (4.17) down to $\eta=n^{-1-\epsilon}$. Setting $\eta_{1}=n^{-1+\xi}$ with $\xi\leq\epsilon/100$ and using that $G^{2}=\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}\eta}G(\mathrm{i}\eta)$, we have
$\Big|\big\langle\mathbf{x},G^{z}(\mathrm{i}\eta_{1})\mathbf{y}\big\rangle-\big\langle\mathbf{x},G^{z}(\mathrm{i}\eta_{0})\mathbf{y}\big\rangle\Big|=$ $\Big|\int_{\eta_{0}}^{\eta_{1}}\big\langle\mathbf{x},\big(G^{z}(\mathrm{i}\eta)\big)^{2}\mathbf{y}\big\rangle\mathrm{d}\eta\Big|$
$\leq$ $\int_{\eta_{0}}^{\eta_{1}}\frac{1}{\eta}\sqrt{\big\langle\mathbf{x},\operatorname{Im}G^{z}(\mathrm{i}\eta)\mathbf{x}\big\rangle\big\langle\mathbf{y},\operatorname{Im}G^{z}(\mathrm{i}\eta)\mathbf{y}\big\rangle}\mathrm{d}\eta$ (4.46)
for any deterministic unit vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$, where we also used the Cauchy-Schwarz inequality and the Ward identity $GG^{*}=\eta^{-1}\operatorname{Im}G$. By a simple spectral decomposition of $H^{z}$, it is easy to check that the function $\eta\big\langle\mathbf{x},\operatorname{Im}G^{z}(\mathrm{i}\eta)\mathbf{x}\big\rangle$ is increasing in $\eta>0$. Hence we obtain
$\Big|\big\langle\mathbf{x},G^{z}(\mathrm{i}\eta_{1})\mathbf{y}\big\rangle-\big\langle\mathbf{x},G^{z}(\mathrm{i}\eta_{0})\mathbf{y}\big\rangle\Big|\leq$ $\sqrt{\big\langle\mathbf{x},\operatorname{Im}G^{z}(\mathrm{i}\eta_{1})\mathbf{x}\big\rangle\big\langle\mathbf{x},\operatorname{Im}G^{z}(\mathrm{i}\eta_{1})\mathbf{x}\big\rangle}\int_{\eta_{0}}^{\eta_{1}}\frac{\eta_{1}}{\eta^{2}}\mathrm{d}\eta\prec n^{2\epsilon}$ (4.47)
where we also used that $\big\langle\mathbf{x},\operatorname{Im}G^{z}(\mathrm{i}\eta_{1})\mathbf{x}\big\rangle\prec 1$ from (4.17) and (4.18) for $\eta_{1}=n^{-1+\xi}$ with $\xi<\epsilon/100$. Again using the local law from (4.17) and (4.18) for $\eta_{1}=n^{-1+\xi}$, $\mathbf{x}=\mathbf{e}_{x}$, and $\mathbf{y}=\mathbf{e}_{y}$, we conclude that, for each fixed $|z|\leq C_{0}$ and $t\geq 0$,
$\max_{x,y\in[2n]}\Big|\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{xy}\Big|\prec n^{2\epsilon},\qquad\mbox{~with~}\quad\eta_{0}=n^{-1-\epsilon}.$
Due to the Lipschitz continuity of $G^{z}_{t}$ in both $t$ and $z$, one can show that the above bound indeed holds true simultaneously for any $t\in[0,t_{0}]$ with $t_{0}=100\log n$ and $|z|\leq C_{0}$, i.e.,
$\sup_{t\in[0,t_{0}]}\sup_{|z|\leq C_{0}}\left\{\max_{x,y\in[2n]}\Big|\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{xy}\Big|\right\}\prec n^{2\epsilon},\qquad\mbox{~with~}\quad\eta_{0}=n^{-1-\epsilon},$ (4.48)
using a grid argument and union bound. Hence using that $\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu}$, we conclude from (4.45) that
$\sup_{t\in[0,t_{0}]}\Big|\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h_{aB}}\Big|\prec n^{2\epsilon+\nu}.$ (4.49)
The same result also applies to $\partial/\partial\overline{h_{aB}}$. In general, for any $k\geq 1$, using the differentiation rules in (4.37) and the local law estimates in (4.48) repeatedly, we obtain the estimate for any $k$-th derivatives of $I^{\geq\eta_{0}}_{t}$ as stated in (4.39). This ends the proof of Lemma 4.4. ∎
5. Lemmas on weak convergence
We collect here classical results used in the proof of Theorem 1.2. The first one states that we can consider compactly supported smooth functions instead of continuous ones in the definition of weak convergence. This is a natural and familiar variant of the Portmanteau theorem that we were not able to locate in a textbook. The second one states that the convergence of point processes is characterized by compactly supported smooth test functions. The third and final one is about the delta method in the context of point process.
Lemma 5.1 (Characterization of convergence in law).
If $X,X_{1},X_{2},\ldots$ are random variables such that
$\lim_{n\to\infty}\mathbb{E}[F(X_{n})]=\mathbb{E}[F(X)]$ (5.1)
for every $\mathcal{C}^{\infty}$ compactly supported $F:\mathbb{R}\to\mathbb{R}$, then the sequence ${(X_{n})}_{n\geq 1}$ converges in law to $X$.
Proof.
First, notice that if (5.1) is true for compactly supported smooth functions, then it is true for compactly supported continuous functions. Indeed, any compactly supported continuous function $F:\mathbb{R}\to\mathbb{R}$ is a uniform limit of compactly supported smooth functions $F_{k}$. Then, for every $k\geq 1$ we have
$\limsup_{n\to\infty}\mathbb{E}[F_{k}(X_{n})]-\|F-F_{k}\|_{\infty}\leq\liminf_{n\to\infty}\mathbb{E}[F(X_{n})]\leq\limsup_{n\to\infty}\mathbb{E}[F(X_{n})]\leq\limsup_{n\to\infty}\mathbb{E}[F_{k}(X_{n})]+\|F-F_{k}\|_{\infty}$
which implies the desired equality by taking $k\to\infty$.
Next, assuming that (5.1) is true for compactly supported smooth functions, we may show that $(X_{n})_{n\geq 1}$ is tight. For this, we may take a smooth function $g:\mathbb{R}\to[0,1]$ supported on [1,1][-1,1] such that $g|_{[-1/2,1/2]}\equiv 1$. Using that $\lim_{M\to\infty}g(x/M)=1$ and that $g$ is bounded by $1$, we obtain that $\lim_{M\to\infty}\mathbb{E}[g(X/M)]=1$. Then, for any $\varepsilon>0$, we may find an $M>0$ such that $\mathbb{E}[f(X/M)]>1-\varepsilon$. Finally, since $\lim_{n\to\infty}\mathbb{E}[g(X_{n}/M)]=\mathbb{E}[g(X/M)]$, we know that $\mathbb{E}[g(X_{n}/M)]>1-\varepsilon$ for $n$ large enough so that
$\mathbb{P}(|X_{n}|\leq M)\geq\mathbb{E}[g(X_{n}/M)]>1-\varepsilon.$
Finally, if $F$ is a bounded continuous function, we may write F(x)=F(x)g(x/M)+F(x)(1g(x/M))F(x)=F(x)g(x/M)+F(x)(1-g(x/M)) so that
$\mathbb{E}[F(X)g(X/M)]-\|F\|_{\infty}\sup_{n\geq 1}\mathbb{P}(|X_{n}|\geq M/2)\leq\liminf_{n\to\infty}\mathbb{E}[F(X_{n})]$
and
$\limsup_{n\to\infty}\mathbb{E}[F(X_{n})]\leq\mathbb{E}[F(X)g(X/M)]+\|F\|_{\infty}\sup_{n\geq 1}\mathbb{P}(|X_{n}|\geq M/2).$
Taking $M\to\infty$ we get the desired equality.
We may have also done a quicker proof by using Prokhorov’s theorem (since we already showed tightness) and noticing that a random variable is characterized by the expected values of smooth compactly supported functions. ∎
Lemma 5.2 (Point processes).
Let $\mathcal{X},\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ be point processes on $\mathbb{C}$ such that
$\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}_{n}}\xrightarrow[n\to\infty]{\mathrm{law}}\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}}$
for every smooth compactly supported function $f:\mathbb{C}\to\mathbb{R}$. Then, the same holds for every continuous compactly supported function $f:\mathbb{C}\to\mathbb{R}$, in other words the sequence $(\mathcal{X}_{n})_{n\geq 1}$ converges in law towards $\mathcal{X}$.
Proof.
Let us show that for all bounded Lipschitz $F:\mathbb{R}\to\mathbb{R}$ and continous and compactly supported $f:\mathbb{C}\to\mathbb{R}$,
$\lim_{n\to\infty}\mathbb{E}\Big[F\Big(\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}_{n}}\Big)\Big]=\mathbb{E}\Big[F\Big(\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}}\Big)\Big].$
Let $(f_{k})_{k\geq 1}$ be a sequence of smooth functions whose support is included in the common compact set $K$ with $\|f-f_{k}\|_{\infty}\to 0$. By denoting $\nu_{n}=\mu_{\mathcal{X}_{n}}$ and $\nu_{\infty}=\mu_{\mathcal{X}}$, we can write
$\Big|\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{n}\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{\infty}\Bigr)\Bigr]\Big|\leq T_{n}^{(k)}+T_{\infty}^{(k)}+S_{n}^{(k)},$
where
$T_{n}^{(k)}=\Big|\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{n}\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f_{k}\mathrm{d}\nu_{n}\Bigr)\Bigr]\Big|\quad\mbox{ and }\quad S^{(k)}_{n}=\Big|\mathbb{E}\Bigr[F\Bigr(\int f_{k}\mathrm{d}\nu_{n}\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f_{k}\mathrm{d}\nu_{\infty}\Bigr)\Bigr]\Big|.$
Since $F$ is Lipschitz and bounded, there is a constant $C>0$ such that $|F(x)-F(y)|\leq C\min\{|x-y|,1\}$. We have
$\limsup_{n\to\infty}T_{n}^{(k)}\leq\limsup_{n\to\infty}\mathbb{E}\big[C\min\big\{\|f-f_{k}\|_{\infty}\nu_{n}(K),1\big\}\big]\leq\limsup_{n\to\infty}\mathbb{E}\Big[H^{(k)}\Big(\int g\mathrm{d}\nu_{n}\Big)\Big],$
where we have taken a compactly supported smooth function $g:\mathbb{C}\to\mathbb{R}$ such that $1_{K}\leq g$ and we have defined $H^{(k)}(x)=C\min\{\|f-f_{k}\|_{\infty}|x|,1\}$. Since $H^{(k)}$ is continuous and bounded we get
$\limsup_{n\to\infty}T_{n}^{(k)}\leq\mathbb{E}\Big[H^{(k)}\Big(\int g\mathrm{d}\nu_{\infty}\Big)\Big]$
so that we have
$\limsup_{n\to\infty}\Big|\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{n}\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{\infty}\Bigr)\Bigr]\Big|\leq 2\mathbb{E}\Big[H^{(k)}\Big(\int g\mathrm{d}\nu_{\infty}\Big)\Big].$
Since $H^{(k)}$ converges pointwise to zero as $k$ goes to $\infty$, we may apply the dominated convergence theorem to $H^{(k)}(\int g\mathrm{d}\nu_{\infty})$ to get $\lim_{k\to\infty}2\mathbb{E}[H^{(k)}(\int g\mathrm{d}\nu_{\infty})]=0$ so that
$\limsup_{n\to\infty}\Big|\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{n}\Bigr)\Bigr]-\mathbb{E}\Bigr[F\Bigr(\int f\mathrm{d}\nu_{\infty}\Bigr)\Bigr]\Big|=0,$
which concludes the proof.
The following lemma is a consequence of the classical fact that if $f_{n}:M\to M$ is a sequence of measurable functions on a locally compact metric space $M$ converging uniformly on compact sets to the identity function and $\mathcal{X}_{n}$ is a sequence of point processes converging to $\mathcal{X}$ then $f_{n}(\mathcal{X}_{n})$ also converges to $\mathcal{X}$. For convenience of the reader, we give a direct proof in the particular case that interests us which can be seen as a “delta method” lemma.
Lemma 5.3 (Delta-method for point processes).
Let $(\mathcal{X}_{n})_{n\geq 1}$ be a sequence of point processes on an open set $U\subset\mathbb{C}$ and fix $x_{0}\in U$. Consider two open sets $V_{1},V_{2}\subset\mathbb{C}$ containing the origin together with diffeomorphisms $\varphi_{i}:U\to V_{i}$. Then $\sqrt{n}(\varphi_{1}(\mathcal{X}_{n})-\varphi_{1}(x_{0}))$ converges in law if and only if $\sqrt{n}(\varphi_{2}(\mathcal{X}_{n})-\varphi_{2}(x_{0}))$ also does and, in that case,
$\lim_{n\to\infty}\sqrt{n}\big(\varphi_{1}(\mathcal{X}_{n})-\varphi_{1}(x_{0})\big)=(\mathrm{d}\varphi_{1})_{x_{0}}\circ(\mathrm{d}\varphi_{2})_{x_{0}}^{-1}\Big(\lim_{n\to\infty}\sqrt{n}\big(\varphi_{2}(\mathcal{X}_{n})-\varphi_{2}(x_{0})\big)\Big).$
Proof.
First, by changing $\varphi_{i}$ by $\varphi_{i}-\varphi_{i}(x_{0})$, we may consider $\varphi_{i}(x_{0})=0$. Then, under this assumption, by considering the point process $\mathcal{Y}_{n}=\varphi_{1}(\mathcal{X}_{n})$ and the maps $\psi_{i}=\varphi_{i}\circ\varphi_{1}^{-1}$ the statement reduces to the following statement. Let $\mathcal{Y}_{n}$ be a sequence of point processes on $V_{1}$ that contains the origin such that $\sqrt{n}\mathcal{Y}_{n}$ converges. If $\psi:V_{1}\to V_{2}$ is a diffeomorphism that satisfies $\psi(0)=0$ then
$\lim_{n\to\infty}\sqrt{n}\psi(\mathcal{Y}_{n})=(\mathrm{d}\psi)_{0}\big(\lim_{n\to\infty}\sqrt{n}\mathcal{Y}_{n}\big).$
To show this, let us consider a compactly supported smooth function $f:\mathbb{C}\to\mathbb{R}$ and let us study the difference $\sum_{y\in\mathcal{Y}_{n}}f(\sqrt{n}\psi(y))-\sum_{y\in\mathcal{Y}_{n}}f(\sqrt{n}\mathrm{d}\psi_{0}(y))$. Write $\psi(x)=\mathrm{d}\psi_{0}(x)+\|x\|r(x)$, where the limit of $r(x)$ as $x$ goes to zero is zero. Since $f$ is smooth and compactly supported it is Lipschitz with some Lipschitz constant $C>0$ and its support is contained in some disk $\mathcal{D}$. Using that $\psi^{-1}$ is Lipschitz near the origin we can take a disk $\widetilde{\mathcal{D}}$ such that $\psi^{-1}(\mathcal{D}/\sqrt{n})\subset\widetilde{\mathcal{D}}/\sqrt{n}$ and we may also assume $\mathrm{d}\psi_{0}^{-1}(\mathcal{D}/\sqrt{n})\subset\widetilde{\mathcal{D}}/\sqrt{n}$. Consider a compactly supported continuous function $g:\mathbb{C}\to\mathbb{R}$ satisfying $1_{\widetilde{\mathcal{D}}}\leq g$. We may write
$\Big|\sum_{y\in\mathcal{Y}_{n}}f(\sqrt{n}\psi(y))-\sum_{y\in\mathcal{Y}_{n}}f(\sqrt{n}\mathrm{d}\psi_{0}(y))\Big|$ $\leq\sum_{y\in\mathcal{Y}_{n}}C\sqrt{n}|\psi(y)-\mathrm{d}\psi_{0}(y)|1_{y\in\psi^{-1}(\mathcal{D}/\sqrt{n})\cup\mathrm{d}\psi_{0}^{-1}(\mathcal{D}/\sqrt{n})}$
$\leq\sum_{y\in\mathcal{Y}_{n}}C\sqrt{n}\|y\||r(y)|1_{\widetilde{\mathcal{D}}/\sqrt{n}}(y).$ (5.2)
If $R$ is the radius of $\widetilde{\mathcal{D}}$, we know that $\sqrt{n}\|y\|1_{\widetilde{\mathcal{D}}/\sqrt{n}}(y)\leq R$. Since $\lim_{x\to 0}r(x)=0$, we know that $r(y)1_{\widetilde{\mathcal{D}}/\sqrt{n}}(y)\leq\varepsilon_{n}$ for a sequence $\varepsilon_{n}$ that goes to zero. Then we can bound (5.2) by
$CR\varepsilon_{n}\sum_{y\in\mathcal{Y}_{n}}g(\sqrt{n}y)\xrightarrow[n\to\infty]{\mathrm{law}}0.$
This implies the result. ∎